OpenStudy (masechaba):

Use De Moivre's Theorem to express cos 6∅ in terms of sums and differences of powers of sin∅ and cos∅

OpenStudy (anonymous):

\[\Large [r(\cos \theta+i\sin\theta)]^n = r^n(\cos n\theta+i\sin n\theta) \]

OpenStudy (sirm3d):

follow @wio 's idea. set \(\large r=1\) \[\large (\cos \theta + i\sin \theta)^n=\cos(n\theta)+i\sin(n\theta)\] expand the left hand side, then compare (and equate) real parts of the complex numbers

OpenStudy (masechaba):

ok will do thanks guys.